4.If 3 workers can paint a room in 2 hours, then approximately how long does it take 4 workers to paint the same room? Assume the time needed to paint the room is inversely proportional to the number of workers.
A. 1 hour
B. 3 hours
C. 1.5 hours
D. 2 hours

Answer by jorel1380(3719)

If 3 workers can paint a room in 2 hours, then one worker can paint 1/3/2 or 1/6 of a room per hour. 1/(4/6)=6/4 hours for 4 people to paint a room= 1-1/2 hours..